OpenStudy (anonymous):

add any three numbers of the following and answer should be 60 2 , 6, 10, 14, 18, 22, 26, 30, 34, 38, 42, 46, 50, 54, 58

OpenStudy (jtvatsim):

What have you tried so far?

OpenStudy (anonymous):

yea tried. but i didnt get

OpenStudy (jtvatsim):

alright, let's start by getting rid of some choices that don't work. :)

OpenStudy (anonymous):

k

OpenStudy (jtvatsim):

we need to pick 3 numbers that add up to 60

OpenStudy (jtvatsim):

so could we ever pick 58? or is that going to be too big?

OpenStudy (jtvatsim):

for example, if I pick 58, then the smallest thing I could choose would be 2, but then I'm already at 60 with just two numbers, but I wanted to use three!

OpenStudy (jtvatsim):

So, we can cross off 58 from the list, it can't ever work.

OpenStudy (jtvatsim):

does that make sense?

OpenStudy (jtvatsim):

of course, there's more, but I want to make sure you understand so far. :)

OpenStudy (anonymous):

from 2 to 58 any three nos. must be added and answer should get 60. my lecturer gave this question. its like puzzle. im confused

OpenStudy (jtvatsim):

OK, I understand, but do you see that we can never use 58 as one of our numbers? it's too big.

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, excellent! So we now only need to think about the numbers from 2 to 54. We've made things easier (just a little). :)

OpenStudy (jtvatsim):

So, let's pretend that I pick 54 as my first number... can I pick a second and third number that will work? Or are we still too big?

OpenStudy (jtvatsim):

54 + ? + ? = 60 Can we choose numbers that make this true?

OpenStudy (jtvatsim):

Well, we can't pick 6 for the second number or else 54 + 6 = 60 TOO BIG!

OpenStudy (jtvatsim):

And 2 doesn't work either, 54 + 2 + 2 = 58 TOO SMALL!

OpenStudy (jtvatsim):

So, 54 is impossible :)

OpenStudy (anonymous):

not same number

OpenStudy (jtvatsim):

Oh, and there you go, so it's even more impossible!

OpenStudy (anonymous):

one number should be used only once

OpenStudy (anonymous):

yes i think so.

OpenStudy (jtvatsim):

so cross off 54, it doesn't work :)

OpenStudy (jtvatsim):

Now, we can keep doing this for each of the numbers.

OpenStudy (jtvatsim):

You will find that 50, 46, and 42 do not work.

OpenStudy (jtvatsim):

And neither does 38.

OpenStudy (jtvatsim):

Let's try 34 as the first number.

OpenStudy (jtvatsim):

If we pick 26 as the second we are too big... 34 + 26 = 60 TOO BIG!

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

So maybe 22? 34 + 22 + ? = 60 means that 56 + ? = 60 ... nope... don't have a 4.

OpenStudy (jtvatsim):

maybe 18? 34 + 18 = 52... nope don't have an 8... AHH!!! A little frustrating right? :)

OpenStudy (jtvatsim):

But, we must be getting closer...

OpenStudy (jtvatsim):

I'm still trying to find the answer too, so the game takes a little bit of time to figure out... :)

OpenStudy (jtvatsim):

hmm... maybe I ruled one out that I shouldn't have...

OpenStudy (anonymous):

im trying it n searching on net. still i didnt get

OpenStudy (jtvatsim):

ok, me too :)

OpenStudy (jtvatsim):

it might be an impossible puzzle... there might not be any combinations that work

OpenStudy (jtvatsim):

actually, yes it is not possible to answer this question.... no combination exists.

OpenStudy (anonymous):

ok

OpenStudy (jtvatsim):

looks like your lecturer either made a mistake or was just trying to strain your brain. :)

OpenStudy (jtvatsim):

you aren't imagining things though, I worked it out with some algebra and it is impossible as far as I can tell. :)

OpenStudy (jtvatsim):

good work though, you persevered on a very difficult question!

OpenStudy (jtvatsim):

I will post the reason it is impossible below, in case you or anyone else cares to know how I solved it. :)

OpenStudy (jtvatsim):

The question wants three different numbers from the list that sum to 60. In other words, X + Y + Z = 60. Ask yourself, what do the numbers on the list look like? They look like 2, 6, 10, 14, etc. all of which can be calculated by the rule 2 + 4*(some whole number). I will call X = 2 + 4*a, Y = 2 + 4*b, and Z = 2 + 4*c. Where a, b, c are whole numbers. Then, substituting we would need (2 + 4*a) + (2 + 4*b) + (2 + 4*c) = 60 6 + 4*a + 4*b + 4*c = 60 4*a + 4*b + 4*c = 54 4*(a + b + c) = 54 so, a + b + c = 54/4. But 54/4 is 13.5 which is NOT a whole number and it should be since a, b, and c are all whole numbers! Thus, the question is ridiculous and impossible to solve.

OpenStudy (anonymous):

may be ur correct. i think my lecturer gave us the brain work.

OpenStudy (jtvatsim):

probably :) it's certainly a good (but very frustrating) exercise.